# Mixed-precision Neural Networks on RISC-V Cores: ISA extensions for Multi-Pumped Soft SIMD Operations

### v1

@@ Abstract.
Recent advancements in quantization and mixed-precision approaches offers substantial opportunities to improve the speed and energy efficiency of Neural Networks (NN). Research has shown that individual parameters with varying low precision, can attain accuracies comparable to full-precision counterparts. However, modern embedded microprocessors provide very limited support for mixed-precision NNs regarding both Instruction Set Architecture (ISA) extensions and their hardware design for efficient execution of mixed-precision operations, i.e., introducing several performance bottlenecks due to numerous instructions for data packing and unpacking, arithmetic unit under-utilizations etc. In this work, we bring together, for the first time, ISA extensions tailored to mixed-precision hardware optimizations, targeting energy-efficient DNN inference on leading RISC-V CPU architectures. To this end, we introduce a hardware-software co-design framework that enables cooperative hardware design, mixed-precision quantization, ISA extensions and inference in cycle-accurate emulations. At hardware level, we firstly expand the ALU unit within our proof-of-concept micro-architecture to support configurable fine-grained mixed-precision arithmetic operations. Subsequently, we implement multi-pumping to minimize execution latency, with an additional soft SIMD optimization applied for 2-bit operations. At the ISA level, three distinct MAC instructions are encoded extending the RISC-V ISA, and exposed up to the compiler level, each corresponding to a different mixed-precision operational mode. Our extensive experimental evaluation over widely used DNNs and datasets, such as CIFAR10 and ImageNet, demonstrates that our framework can achieve, on average, 15$\times$ energy reduction for less than 1% accuracy loss and outperforms the ISA-agnostic state-of-the-art RISC-V cores.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
In recent years, numerous studies have explored various compression optimization techniques, such as pruning and quantization (zervtc, ) in order to minimize the inference latency and memory footprint of DNNs on edge devices. Optimized software libraries such as ARM’s CMSIS-NN (CMSIS-NN, ), Google’s GEMMLowp (jacob2022gemmlowp, ), and X-CUBE-AI from STMicroelectronics (STM-XCUBE-AI, ), along with advanced frameworks like DORY (Dory, ) that targets RISC-V PULP platform processors, and MCUNet (MCUNet, ; CMix-NN, ) are widely used in order to deploy Quantized Neural Networks on commercial devices. Nonetheless, these tools primarily support 8-bit or higher precision variables and therefore cannot fully exploit the computational efficiencies offered by lower-bit quantization of DNNs.
@@
Since most modern day systems are unable to efficiently compute sub-byte calculations, many works have focused on developing specialized DNN accelerators (YodaNN, ; tcad20, ; Eyeriss, ; Barvinn, ; DNN_Intel_FPGA, ). These accelerators often provide significant performance enhancements, but those benefits are often compromised due to their lack of scalability, extensive area occupation, and high power consumption, typically in the range of hundreds of milliwatts (YodaNN, ; tcad20, ; Eyeriss, ) to a few watts (Barvinn, ). These drawbacks render them less suitable for ultra-low power applications where efficiency and compactness are paramount.
@@
As an alternative approach, other works have looked into extending the RISC-V ISA and developing custom functional units specifically designed to handle efficient sub-byte operations. PULP-NN (Pulp-NN, ) achieves significant speed improvements by enabling lower-bit DNN inference, utilizing instructions that pack and extract vectors of smaller data sizes, such as 4-bit and 2-bit. However, these specialized casting instructions introduce overheads for computations at 4-bit and 2-bit sizes, thus diminishing the performance benefits associated with these lower bit widths. On the other hand, its extension XpulpNN (XpulpNN, ) supports 2, 4, and 8-bit SIMD operations but it is not supporting mixed-precision computations.
@@
Recent studies (ottavi2020mixed, ; lee2018unpu, ; Mix-GEMM, ; SySMOL, ) examined the utilization of mixed precision variables on resource constrained devices, as an effective strategy to balance the trade-offs between accuracy, latency, and memory footprint (wang2024optimizing, ) that arise when employing lower precision numbers. UNPU (lee2018unpu, ) explored bit-serial MAC units with consistent activation data size and weight data sizes spanning 1 to 16 bits, while (ottavi2020mixed, ) extended a RISC-V core by incorporating 4- and 2-bit MAC units alongside customized controllers to enable 2-/4-/8-bit mixed-precision computations. In Mix-GEMM (Mix-GEMM, ) and SySMOL (SySMOL, ), the authors presented software/hardware co-design workflows that seamlessly integrate hardware accelerators into the processor architecture complemented by hardware-aware training. This approach, enhanced by systematic exploration of their respective design space, enables the possibility of fine-grained mixed-precision in DNNs. However, in (Mix-GEMM, ) they only consider per-network quantization in their evaluation, potentially overlooking the benefits of more granular precision adjustments, while (SySMOL, ) restricts its use to 1, 2, and 4-bit variables within intra-layer quantization, which can lead to significant accuracy losses in more complex models or challenging datasets like ImageNet. Furthermore, neither study explores low-level hardware optimizations, such as overclocking (Approximate_Computing_Survey, ) or packing multiple low-precision multiplications on a single block (FILMQNNEF, ) to optimize their system’s throughput.
@@
Table 1 summarizes the above discussion and provides a qualitative comparison of most relevant works in the field of RISC-V ISA extension for mixed-precision DNNs. Our work distinguishes from most state-of-the-art approaches, since, to the best of our knowledge, none has proposed mixed-precision instructions capable of simultaneously activating different computations and operational modes through logic and circuit-level optimizations, such as multi-pumping and soft SIMD techniques. In Section 5.3 we thoroughly evaluate our work against most relevant state-of-the-art approaches.
@@ 3. Configurable mixed-precision architecture based on Ibex core
The base design utilized in this work is a generic microarchitecture implementation of the Ibex (ibexproc, ), an open-source 32 bit RISC-V CPU core, depicted in Fig. 1. Without loss of generality, the specific RISC-V core forms a proof-of-concept micro-architecture scenario to showcase the impact when enabling in an execution unit the support of fined-grained mixed-precision operations and can be also generalized in a straightforward manner to other RISC-V CPU cores.
@@ 3.3. RISC-V ISA Extension
To put the speedup of our modes into perspective, we extend RISC-V ISA to support three different operations. Our encoding follows the rules of the RISC-V ISA manual for custom extensions (waterman2014risc, ). The three instructions adhere to the R-type structured format comprising six sub-fields, i.e., opcode, func3, func7, rs1, rs2 and rd, and has a fixed length of 32 bits. The opcode dictates the operation to be executed and the involved operands, in both destination (rd) and source (rs1,rs2) registers. Func3 and func7 describe the operations depending on the input format given to them. Our instruction listing for the RISC-V mixed-precision extension is shown in Table 2. As shown, we introduce only a few instructions, having thus zero impact on the hardware of the CPU’s opcode decoder. Specifically, the $nn\_mac\_8b$ corresponds to $Mode-1$ and requires 4 packed weights of 8-bit to perform 4 MAC operations in parallel, while $nn\_mac\_4b$ and $nn\_mac\_2b$ correspond to $Mode-2$ and $Mode-3$, respectively, and require 8 and 16 packed weights, of 4- and 2-bits. Note, that in all cases the length of the accumulator (rd) is 32-bit. Once all the required MAC operations for each output feature are completed, the final step involves fetching the results from the 32-bit accumulators. Then, a common requantization step (jacob2018quantization, ) is performed to adjust the bit-width of the outputs back to 8-bits.

### v2

@@ Abstract.
Recent advancements in quantization and mixed-precision approaches offers substantial opportunities to improve the speed and energy efficiency of Neural Networks (NN). Research has shown that individual parameters with varying low precision, can attain accuracies comparable to full-precision counterparts. However, modern embedded microprocessors provide very limited support for mixed-precision NNs regarding both Instruction Set Architecture (ISA) extensions and their hardware design for efficient execution of mixed-precision operations, i.e., introducing several performance bottlenecks due to numerous instructions for data packing and unpacking, arithmetic unit under-utilizations etc. In this work, we bring together, for the first time, ISA extensions tailored to mixed-precision hardware optimizations, targeting energy-efficient DNN inference on leading RISC-V CPU architectures. We introduce a hardware-software co-design framework that supports cooperative hardware design, mixed-precision quantization, ISA extensions, and cycle-accurate emulations. At the hardware level, we expand the ALU unit in our micro-architecture for configurable mixed-precision arithmetic operations and implement multi-pumping to reduce execution latency, with soft SIMD optimization for 2-bit operations. At the ISA level, we encode three distinct MAC instructions extending the RISC-V ISA, each for different mixed-precision modes, and expose them to the compiler. Our extensive experimental evaluation over widely used DNNs and datasets, such as CIFAR10 and ImageNet, demonstrates that our framework can achieve, on average, 15$\times$ energy reduction for less than 1% accuracy loss and outperforms the ISA-agnostic state-of-the-art RISC-V cores.
@@
journalyear: 2024copyright: rightsretainedconference: IEEE/ACM International Conference on Computer-Aided Design; October 27–31, 2024; New York, NY, USAbooktitle: IEEE/ACM International Conference on Computer-Aided Design (ICCAD ’24), October 27–31, 2024, New York, NY, USAdoi: 10.1145/3676536.3676840isbn: 979-8-4007-1077-3/24/10
@@ 1. Introduction
In recent years, numerous studies have explored various compression optimization techniques, such as pruning and quantization (zervtc, ), to minimize the inference latency and memory footprint of DNNs on edge devices. Optimized software libraries such as ARM’s CMSIS-NN (CMSIS-NN, ), Google’s GEMMLowp (jacob2022gemmlowp, ), and X-CUBE-AI from STMicroelectronics (STM-XCUBE-AI, ), along with advanced frameworks like DORY (Dory, ) for RISC-V PULP platform processors and MCUNet (MCUNet, ; CMix-NN, ), are widely used to deploy Quantized Neural Networks on commercial devices. Nonetheless, these tools primarily support 8-bit or higher precision variables and cannot fully exploit the computational efficiencies offered by lower-bit quantization of DNNs.
@@
Since most modern systems are unable to efficiently compute sub-byte calculations, many works have focused on developing specialized DNN accelerators (YodaNN, ; tcad20, ; Eyeriss, ; Barvinn, ; DNN_Intel_FPGA, ). These accelerators often provide significant performance enhancements, but those benefits are often compromised due to their lack of scalability, extensive area occupation, and high power consumption, typically in the range of hundreds of milliwatts (YodaNN, ; tcad20, ; Eyeriss, ) to a few watts (Barvinn, ).
@@
Alternatively, other works have looked into extending the RISC-V ISA and developing custom functional units specifically designed to handle efficient sub-byte operations. PULP-NN (Pulp-NN, ) achieves significant speed improvements by enabling lower-bit DNN inference, utilizing instructions that pack and extract vectors of smaller data sizes, such as 4-bit and 2-bit. However, these specialized casting instructions introduce overheads for computations at 4-bit and 2-bit sizes, diminishing the performance benefits associated with these lower bit-widths. XpulpNN (XpulpNN, ) supports 2, 4, and 8-bit SIMD operations but does not support mixed-precision computations.
@@
Recent studies (ottavi2020mixed, ; lee2018unpu, ; Mix-GEMM, ; SySMOL, ) examined the utilization of mixed precision variables on resource-constrained devices as an effective strategy to balance the trade-offs between accuracy, latency, and memory footprint (wang2024optimizing, ) that arise when employing lower precision numbers. UNPU (lee2018unpu, ) explored bit-serial MAC units with consistent activation data size and weight data sizes spanning 1 to 16 bits, while (ottavi2020mixed, ) extended a RISC-V core by incorporating 4- and 2-bit MAC units alongside customized controllers to enable 2-/4-/8-bit mixed-precision computations. In Mix-GEMM (Mix-GEMM, ) and SySMOL (SySMOL, ), the authors presented software/hardware co-design workflows that seamlessly integrate hardware accelerators into the processor architecture complemented by hardware-aware training. However, in (Mix-GEMM, ) they only consider per-network quantization in their evaluation, potentially overlooking the benefits of more granular precision adjustments, while (SySMOL, ) restricts its use to 1, 2, and 4-bit variables within intra-layer quantization, which can lead to significant accuracy losses in more complex models or challenging datasets like ImageNet. Furthermore, neither study explores low-level hardware optimizations, such as overclocking (Approximate_Computing_Survey, ) or packing multiple low-precision multiplications on a single block (FILMQNNEF, ) to optimize their system’s throughput.
@@
Table 1 summarizes the above discussion and provides a qualitative comparison of most relevant works in the field of RISC-V ISA extension for mixed-precision DNNs. Our work distinguishes from most state-of-the-art approaches, since, to the best of our knowledge, none has proposed mixed-precision instructions capable of simultaneously activating different computations and operational modes through logic and circuit-level optimizations, such as multi-pumping and soft SIMD techniques. In Section 5.3 we evaluate our work against most relevant state-of-the-art approaches.
@@ 3. Configurable mixed-precision architecture based on Ibex core
The base design utilized in this work is a generic microarchitecture implementation of the Ibex (ibexproc, ), an open-source 32 bit RISC-V CPU core (Fig. 1). Without loss of generality, the specific RISC-V core forms a proof-of-concept micro-architecture scenario to showcase the impact when enabling in an execution unit the support of fined-grained mixed-precision operations and can be also generalized in a straightforward manner to other RISC-V CPU cores.
@@ 3.3. RISC-V ISA Extension
To illustrate the speedup of our modes, we extend the RISC-V ISA to support three different operations, following the RISC-V ISA manual for custom extensions. These three instructions use the R-type format with six sub-fields: opcode, func3, func7, rs1, rs2, and rd, all 32 bits long. The opcode specifies the operation and the involved source (rs1, rs2) and destination (rd) registers. Func3 and func7 define the operations based on the input format.
Our RISC-V mixed-precision extension instructions are listed in Table 2. We introduce only a few instructions, having no impact on the CPU’s opcode decoder hardware. Specifically, $nn/_{m}ac/_{8}b$ corresponds to $Mode-1$ and requires four 8-bit packed weights for four parallel MAC operations. $nn/_{m}ac/_{4}b$ and $nn/_{m}ac/_{2}b$ correspond to $Mode-2$ and $Mode-3$, requiring eight 4-bit and 16 2-bit packed weights, respectively. In all cases, the accumulator (rd) length is 32 bits. After completing the MAC operations for each output feature, the results are fetched from the 32-bit accumulators, followed by a common requantization step to adjust the output bit-width back to 8 bits.